# Trees of Graphs as Boundaries of Hyperbolic Groups

### v1

@@ 7. $V$-trees of graphs and their decompositions into core and arms
We present here some alternative expression of the $\mathcal{V}$-tree of $\theta$-graphs ${\mathcal{X}}_{V}[{\mathcal{V}}_{\Xi_{k}}]$ as tree of arms. This expression, a bit simpler than that provided by the tree system $\Psi[{\mathcal{V}}_{\Xi_{k}}]$, will be used in the proof of the implication (3$\Rightarrow$(2) in Theorem A, given in Section 13.
Consider the tree system $\Psi_{w}[{\mathcal{V}}_{\Xi_{k}}]$ described as follows. The underlying tree $T_{w}$ of $\Psi_{w}[{\mathcal{V}}_{\Xi_{k}}]$ consists of a single white vertex $u_{c}$ and precisely $k$ black vertices $t_{A}$ corresponding to the arms $A$ of ${\mathcal{X}}_{V}({\mathcal{V}}_{\Xi_{k}})$, and of precisely $k$ edges of the form $[u_{c},t_{A}]$. As the constituent spaces for $\Psi_{w}[{\mathcal{V}}_{\Xi_{k}}]$ we take $K_{t_{A}}=A$, and as the peripheral space we take $\Sigma_{u_{c}}=C[{\mathcal{V}}_{\Xi_{k}}]$. As the maps $\varphi_{[u_{c},t_{A}]}:\Sigma_{u_{c}}\to K_{t_{A}}$ we take the natural inclusions of form $C[{\mathcal{V}}_{\Xi_{k}}]\subset A$.
We skip a straightforward argument for showing that the inclusions $A\subset{\mathcal{X}}_{V}({\mathcal{V}}_{\Xi_{k}})$ induce a natural identification of the limit space $\lim\Psi_{w}[{\mathcal{V}}_{\Xi_{k}}]$ with the $V$-tree ${\mathcal{X}}_{V}({\mathcal{V}}_{\Xi_{k}})$.
@@
A line $L$ in a V-tree system $\Theta[{\mathcal{V}}]$ is orientable if any two appearances $e_{k}^{\circ}$, $e_{m}^{\circ}$ in $L$ of any edge $e$ of $\Gamma$ are consistently oriented. The latter means that in the corresponding arm $A_{L}$, viewed as an internally punctured segment, the orientations of subsegments $e_{k}^{\circ}$ and $e_{m}^{\circ}$ induced from a fixed orientation of $e$ are consistent. A line $L$ is non-orientable, if some edge $e$ of $\Gamma$ has some two inconsistently oriented appearances in $L$. Accordingly with that, we speak of orientability and non-orientability of the corresponding arms $A_{L}$.
@@
[…]
@@
To prove the converse implication, observe that all appearances of any oriented edge $\varepsilon$ of $\Gamma$ in lines of $\Theta[{\mathcal{V}}]$ are followed by adjacent appearances of some uniquely determined by $\varepsilon$ other oriented edge (which is not necessarily distinct from $\varepsilon$). Thus, if some line contains oppositely oriented appearances of some edge, it must also contain, somewhere in between, oppositely oriented and adjacent appearances of some other edge, which clearly implies existence of $v$ and $p$ as in the statement. This completes the proof. ∎
@@
    If all lines in ${\mathcal{L}}_{P}$ are orientable then any fixed order of the doubleton $P$ induces orientations on all edges from $E_{\Gamma}^{P}$. We again call the so described orientations the order-induced orientations (with respect to the above fixed order of the peripheral $P$).
@@ 8.1. Abstract factors and their decorations
It is obviously true that the conjugate subgroup $gH_{i}g^{-1}$ depends only on the coset in $K/H_{i}$ to which the conjugating element $g$ belongs; for this reason we will speak of conjugates $gH_{i}g^{-1}$ by the cosets $g\in K/H_{i}$. Since it is well known that for distinct maximal 2-ended subgroups $H,H^{\prime}$ of a hyperbolic group $K$ their boundaries $\partial H,\partial H^{\prime}$ viewed as subsets in the Gromov boundary $\partial K$ are disjoint, it follows that the family of boundaries of the conjugates
@@
consists of pairwise disjoint subsets of $\partial K$. It is also not hard to observe that this family of subsets of $\partial K$ is null. To see it, note that any conjugate $gH_{i}g^{-1}$ is at finite Hausdorff distance (for any word metric $d_{K}$ in $K$) from the coset $gH$, and thus we have $\partial(gHg^{-1})=\partial(gH)$. Moreover, for any fixed $i\in I$ and for the identity element $e\in K$ the distances
@@
clearly diverge to infinity. Since the cosets $gH_{i}$ of $G/H_{i}$ are uniformly quasiconvex in $G$, it follows that the diameters $\hbox{diam}(\partial(gH_{i})):gH_{i}\in G/H_{i}$ converge to 00, which obviously implies that the family (1) is null. This allows to introduce the following concept.
@@
$(\partial K,\{\partial(gH_{i}g^{-1}):i\in I,g\in K/H_{i}\}).$
Peripherals of form $\partial(gH_{i}g^{-1})$ are called $H_{i}$-peripherals, or peripherals of type $H_{i}$.
@@
A maximal 2-ended subgroup $H<K$ of a hyperbolic group $K$ is orientable if each element $h\in H$ induces a homeomorphism of $\partial K$ which fixes the subset $\partial H$ pointwise. Otherwise $H$ is non-orientable. If $H$ is orientable, an orientation for $H$ is a choice of any order in the set $\partial H$.
Note that $H$ as above is non-orientable iff for some $h\in H$ the induced homeomorphism of $\partial K$ transposes the points of the doubleton $\partial H$. Obviously, if $H$ is orientable then any conjugate of $H$ in $K$ is also orientable, and a similar property holds for non-orientability. Moreover, if $H$ is orientable, any fixed orientation for $H$ induces orientations for all conjugates $gHg^{-1}$. (Here we use the fact that, since $K$ is hyperbolic, the stabilizer of the doubleton $\partial H\subset\partial K$ is a 2-ended group containing $H$, so it coincides with $H$, because $H$ is maximal 2-ended.) In particular, an orientation for $H$ induces also orders on the boundary sets $\partial(gHg^{-1})$ (which are doubletons).
@@
An orientation data for an abstract factor $(K,\{H_{i}\}_{i\in I})$ is an information about orientability and non-orientability of the subgroups $H_{i}$ (and thus of all subgroups from their conjugacy classes) and, for each orientable subgroup $H_{i}$, a choice of an orientation. Such an orientation data induces a decoration of the associated space with peripherals, which consists of informations about the types ($H_{i}$-types) of the peripherals, and about their orientability, and if a peripheral is orientable, this decoration indicates one of the orientations of this peripheral, compatible with the fixed orientation of the corresponding 2-ended subgroup.
@@
    Each homeomorphism of the action of $K$ on $\partial K$ preserves any decoration of the associated space with peripherals. More precisely, this means that any such homeomorphism preserves the types of the peripherals, it maps non-orientable peripherals to non-orientable ones, and it maps orientable peripherals to orientable ones, and in the latter case, it preserves the orientations indicated by a decoration.
@@ 8.2. Flexible factors
An abstract flexible factor is an abstract factor $(K,\{H_{i}\}_{i\in I})$ such that the associated space with peripherals is homeomorphic (as a space with peripherals) to the punctured circle $(S^{1})^{\circ}$.
@@
Let $G(v)$ be a flexible factor of the Bowditch JSJ-splitting of a 1-ended hyperbolic group $G$. Then $(G(v),{\mathcal{H}}_{v})$, i.e. the abstract factor associated to $G(v)$ as in Example 8.2, is an abstract flexible factor, in the sense of Definition 8.7. For justification, see Section 5 in [1], where flexible factors are called maximal hanging Fuchsian (MHF), or the paragraph after Definition 5.9 in [15].
@@
    Recall that the natural embedding of the punctured circle $(S^{1})^{\circ}$ in the circle $S^{1}$ is uniquely described, up to a homeomorphism of the ambient $S^{1}$, by the following condition: for any peripheral doubleton $P$ of $(S^{1})^{\circ}$ there is a segment $I_{P}$ contained in $S^{1}$ whose set of endpoints coincides with $P$ and whose interior is disjoint with $(S^{1})^{\circ}$. We will call this embedding the natural extension of $(S^{1})^{\circ}$ to a circle. Each of the segmants $I_{P}$ as above will be called the segment induced by $P$. Note that any homeomorphism $h$ of a space with peripherals $(S^{1})^{\circ}$ (i.e. a homeomorphism which preserves the family of peripherals) can be extended in an essentially unique way to a homeomorphism $\bar{h}:S^{1}\to S^{1}$ of the natural extension. Thus, it makes sense to speak of orientation-preserving and orientation-reversing homeomorphisms $h$, by referring to the corresponding property of the induced $\bar{h}$.
@@
Note that an abstract flexible factor $(K,\{H_{i}\}_{i\in I})$ is non-orientable if for some $g\in K$ the induced homeomorphism of $\partial K=(S^{1})^{\circ}$ is orientation-reversing. Next lemma relates orientability properties of an abstract flexible factor $(K,\{H_{i}\}_{i\in I})$ with orientability properties of its peripheral subgroups $H_{i}$. We omit its straightforward proof.
@@
The reader should keep in mind that, given an abstract flexible factor $\mathcal{F}$, orientability of its all peripheral subgroups does not necessarily lead to orientability of the factor. For example, if $K$ is the fundamental group of a compact connected hyperbolic non-orientable surface with non-empty boundary, and if we take as its peripherals the subgroups corresponding to boundary components, then the obtained flexible factor is non-orientable, though its all peripheral subgroups are cyclic, and hence orientable.
@@
Given an orientable abstract flexible factor ${\mathcal{F}}=(K,\{H_{i}\}_{i\in I})$, with fixed one of the orientations, and an orientable peripheral subgroup $H\in\{H_{i}\}_{i\in I}$, also with fixed one of the orientations (i.e. an order of the doubleton $P=\partial H$), we say that these orientations are compatible if the induced from $P$ orientation of the induced segment $I_{P}$ in the extended circle $S^{1}$ coincides with the orientation of this segment induced from the orientation of the whole $S^{1}$ which represents the fixed orientation of $\mathcal{F}$.
@@
    Suppose that $\mathcal{F}$ is non-orientable. Consider the extended circle $S^{1}$ for $\partial K=(S^{1})^{\circ}$, and fix one of its orientations. If a subgroup $H_{i}$ is orientable, and if one of its orientations is fixed, then the family of those $H_{i}$-peripherals whose orientations are compatible (respectively, incompatible) with the above fixed orientation of $S^{1}$ is non-empty, and the union of this family is dense in $\partial K=(S^{1})^{\circ}$.
@@ 8.3. Abstract virtually free factors and their Whitehead graphs
In this subsection we first recall the concept of a Whitehead graph, as applied to abstract factors that are free. We then extend all of our discussion to abstract virtually free factors. The main result in this subsection, Corollary 8.25, identifies the space with peripherals associated to an abstract virtually free factor with the core of the associated V-tree of Whitehead graphs. This result will play a crucial role in our proof of the implication (3$\Rightarrow$(2) of Theorem A, in Section 13, and it will be applied to virtually free rigid cluster factors of the reduced Bowditch JSJ splitting of a group $G$.
@@
To each abstract free factor $(F,{\mathcal{H}})$, equipped with a basis $B$ of free generators for $F$, there is associated some finite graph, called the Whietehead graph of the factor, see Seection 2E in [4]. We now recall the description of this graph.
@@
Given an abstract free factor ${\mathcal{Q}}=(F,\{H_{i}\})_{1\leq i\leq k}$, let $B=(x_{1},\dots,x_{m})$ be some base set of free generators for $F$. For each $1\leq i\leq k$ fix a generator $h_{i}$ of the cyclic subgroup $H_{i}$, and denote by $w_{i}$ the unique cyclically reduced word over $B\cup B^{-1}$ in the conjugacy class of $h_{i}$. The Whitehead graph for $\mathcal{Q}$ with respect to $B$ is the graph $W=W_{\mathcal{Q}}$ with the vertex set $V_{W_{\mathcal{Q}}}=\{x_{1},\dots,x_{m},x_{1}^{-1},\dots,x_{m}^{-1}\}$ and with one edge connecting distinct vertices $x,y\in V_{W_{\mathcal{Q}}}$ for every occurence of the subword $x^{-1}y$ in the cyclic words $w_{1},\dots,w_{k}$. (Here we use the convention that if $w_{i}=x$ has length 1 then there is precisely one occurence of $xx$ in $w_{i}$, and no occurence of any other word of length 2; moreover, if $w_{i}=uz$ has length 2, then $w_{i}$ has two non-cyclic subwords of length 2, namely $uz$ and $zu$.)
@@
    Observe that, since the subgroups $H_{i}<F$ are maximal cyclic, their generators are not the powers in $F$, and hence the cyclic words $w_{i}$ are not periodic. For this reason, any two occurences of the same subword in any $w_{i}$ are essentially distinct. This dispels some potential doubts concerning the above definition.
@@
    Note that the Whitehead graph $W_{\mathcal{Q}}$ is obviously a graph with no loop edges, but it may contain multiple edges. $W_{\mathcal{Q}}$ depends not only on the factor $\mathcal{Q}$, but also on the choice of a basis $B$ in the free group $F$. The choices for which the number of edges in corresponding Whitehead graphs is minimal lead to so called minimal Whitehead graphs for the factor $\mathcal{Q}$. As we will show later (see Proposition 9.3), in the situations considered in this paper the minimal Whitehead graphs will be always 2-connected (i.e. connected and without cut vertices), so in particular they will be essential (i.e. with no isolated vertices). Since our all further observations concern any choice of a minimal Whitehead graph for a factor $\mathcal{Q}$, we will not bother about this choice, and we will speak of the Whitehead graph associated to $\mathcal{Q}$, meaning any of the minimal Whitehead graphs as above.
Any Witehead graph $W_{\mathcal{Q}}$ which is essential is canonically equipped with some $V$-involution ${\mathbf{a}}_{\mathcal{Q}}$, so that it yields an associated connecting $V$-system ${\mathcal{V}}_{\mathcal{Q}}=(W_{\mathcal{Q}},{\mathbf{a}}_{\mathcal{Q}})$. This $V$-involution consists of the involution $a_{\mathcal{Q}}:V_{W_{\mathcal{Q}}}\to V_{W_{\mathcal{Q}}}$ given by $a_{\mathcal{Q}}(x)=x^{-1}$ for each $x\in V_{W_{\mathcal{Q}}}$ (where we use the convention that $(x_{i}^{-1})^{-1}=x_{i}$), and of the maps $\alpha_{x}:\hbox{Lk}(x)\to\hbox{Lk}(x^{-1})$ described as follows. For any $y\in V_{W_{\mathcal{Q}}}$ and any occurence of $y$ in words $w_{1},\dots,w_{k}$ consider the subwords of form $xy$ and $yz$, one for each of the two forms, which extend this occurence of $y$, and denote by $\sigma,\tau$ the corresponding occurences of these subwords, respectively. Denote by $e_{\sigma}$ and $e_{\tau}$ the edges of $W$ corresponding to the above occurences of $xy$ and $yz$. Denote also by $[e_{\sigma}]_{y}$ and $[e_{\tau}]_{y^{-1}}$ the points in the links $\hbox{Lk}(y)$ and $\hbox{Lk}(y^{-1})$ induced in those links by the edges $e_{\sigma}$ and $e_{\tau}$, respectively. Finally, put $\alpha_{y}([e_{\sigma}]_{y})=[e_{\tau}]_{y^{-1}}$ and $\alpha_{y^{-1}}([e_{\tau}]_{y^{-1}})=[e_{\sigma}]_{y}$; note that the collection of all assignments as above fully describes all the maps $\alpha_{y}:y\in V_{W_{\mathcal{Q}}}$, and that the conditions required in Definition 7.1 are satisfied.
The next observation follows easily from Lemma 7.15, in view of the form of the involution $a_{\mathcal{Q}}$ given above.

### v2

@@ 7. $V$-trees of graphs and their decompositions into core and arms
We present here some alternative expression of the $\mathcal{V}$-tree of $\theta$-graphs ${\mathcal{X}}_{V}[{\mathcal{V}}_{\Xi_{k}}]$ as tree of arms. This expression, a bit simpler than that provided by the tree system $\Psi[{\mathcal{V}}_{\Xi_{k}}]$, will be used in the proof of the implication (3$\Rightarrow$(2) of Theorem A given in Section 13.
Consider the tree system $\Psi_{w}[{\mathcal{V}}_{\Xi_{k}}]$ described as follows. The underlying tree $T_{w}$ of $\Psi_{w}[{\mathcal{V}}_{\Xi_{k}}]$ consists of a single white vertex $u_{c}$ and precisely $k$ black vertices $t_{A}$ corresponding to the arms $A$ of ${\mathcal{X}}_{V}({\mathcal{V}}_{\Xi_{k}})$, and of precisely $k$ edges of the form $[u_{c},t_{A}]$. As the constituent spaces for $\Psi_{w}[{\mathcal{V}}_{\Xi_{k}}]$ we take $K_{t_{A}}=A$, and as the peripheral space we take $\Sigma_{u_{c}}=C[{\mathcal{V}}_{\Xi_{k}}]$. As the maps $\varphi_{[u_{c},t_{A}]}:\Sigma_{u_{c}}\to K_{t_{A}}$ we take the natural inclusions of the form $C[{\mathcal{V}}_{\Xi_{k}}]\subset A$.
Since, by an easy observation, we have a natural identification $\lim\Psi_{w}[{\mathcal{V}}_{\Xi_{k}}]=\lim\Psi[{\mathcal{V}}_{\Xi_{k}}]$, it follows then from Lemma 7.12 that the inclusions $A\subset{\mathcal{X}}_{V}({\mathcal{V}}_{\Xi_{k}})$ induce a natural identification of the limit space $\lim\Psi_{w}[{\mathcal{V}}_{\Xi_{k}}]$ with the $V$-tree ${\mathcal{X}}_{V}({\mathcal{V}}_{\Xi_{k}})$.
@@
A line $L$ in a V-tree system $\Theta[{\mathcal{V}}]$ is orientable if any two appearances $e_{k}^{\circ}$, $e_{m}^{\circ}$ in $L$ of any edge $e$ of $\Gamma$ are consistently oriented. The latter means that in the corresponding arm $A_{L}$, viewed as an internally punctured segment, the orientations of subsegments $e_{k}^{\circ}$ and $e_{m}^{\circ}$ induced from a fixed orientation of $e$ are consistent. A line $L$ is non-orientable, if some edge $e$ of $\Gamma$ has some two inconsistently oriented appearances in $L$. Accordingly, we speak of orientability and non-orientability of the corresponding arms $A_{L}$.
@@
Let ${\mathcal{V}}=(\Gamma,{\mathbf{a}})$ be a connecting V-system. Then the V-tree system $\Theta[{\mathcal{V}}]$ contains a nonorientable line if and only if there is $v\in V_{\Gamma}$ such that $a(v)=v$ and there is $p\in\hbox{Lk}(v)$ such that $\alpha_{v}(p)=p$. In particular, if the involution $a:V_{\Gamma}\to V_{\Gamma}$ is fixed point free then all lines in the V-tree system $\Theta[{\mathcal{V}}]$ are orientable.
@@
To prove the converse implication, observe that all appearances of any oriented edge $\varepsilon$ of $\Gamma$ in lines of $\Theta[{\mathcal{V}}]$ are followed by adjacent appearances of some other oriented edge uniquely determined by $\varepsilon$ (which is not necessarily distinct from $\varepsilon$). Thus, if some line contains oppositely oriented appearances of some edge, it must also contain, somewhere in between, oppositely oriented and adjacent appearances of some other edge, which clearly implies existence of $v$ and $p$ as in the statement. This completes the proof. ∎
@@
    Suppose that the $V$-involution map $a:V_{\Gamma}\to V_{\Gamma}$ of $\mathcal{V}$ is fixed point free. Let $P$ be some doubleton peripheral as in (2) above, and suppose that an edge $e$ of $\Gamma$ has appearances in two distinct lines $L_{1},L_{2}\in\mathcal{L}_{P}$. Fix some order of $P$. The argument similar as in the proof of Lemma 7.15 shows that the order-induced (from the above fixed order of $P$) orientations of $e$ determined by the two above mentioned appearances of $e$ in $L_{1}$ and $L_{2}$ coincide. It follows that any fixed order of the doubleton $P$ uniquely induces orientations on all edges from the set $E_{\Gamma}^{P}$. We again call the so described orientations the order-induced orientations (with respect to the above fixed order of the peripheral $P$).
@@ 8.1. Abstract factors and their decorations
It is obviously true that the conjugate subgroup $gH_{i}g^{-1}<K$ depends only on the coset in $K/H_{i}$ to which the conjugating element $g$ belongs. Moreover, since $H_{i}$ is maximal 2-ended, it coincides with the stabilizer of $\partial H_{i}$ in $\partial G$ (the stabilizer is a 2-ended subgroup containing $H_{i}$) which implies that $H_{i}$ is self-normalizing. Thus the cosets and conjugates of $H_{i}$ are in bijective correspondence and we will freely refer to a conjugate $gH_{i}g^{-1}$ by way of its corresponding coset $gH_{i}\in K/H_{i}$. Since it is well known that for distinct maximal 2-ended subgroups $H,H^{\prime}$ of a hyperbolic group $K$ their boundaries $\partial H,\partial H^{\prime}$ viewed as subsets in the Gromov boundary $\partial K$ are disjoint (see e.g. the sentence right before Lemma 1.1 in [Bow98]), it follows that the family of boundaries of the conjugates
@@
consists of pairwise disjoint subsets of $\partial K$. It is also not hard to observe that this family of subsets of $\partial K$ is null. To see it, note that any conjugate $gH_{i}g^{-1}$ lies at finite Hausdorff distance (for any word metric $d_{K}$ in $K$) from the coset $gH_{i}$, and thus we have $\partial(gH_{i}g^{-1})=\partial(gH_{i})$. Moreover, for any fixed $i\in I$ and for the identity element $e\in K$ the distances
@@
clearly diverge to infinity. Since the cosets $gH_{i}$ of $K/H_{i}$ are uniformly quasiconvex in $K$ (in particular, they lie at uniformly finite distance from geodesics in $K$ connecting their boundary points), it follows that the diameters $\hbox{diam}(\partial(gH_{i})):gH_{i}\in K/H_{i}$ converge to 00, which obviously implies that the family (1) is null. This allows us to introduce the following concept.
@@
$(\partial K,\{\partial(gH_{i}g^{-1}):i\in I,gH_{i}\in K/H_{i}\}).$
Peripherals of the form $\partial(gH_{i}g^{-1})$ are called $H_{i}$-peripherals, or peripherals of type $H_{i}$.
@@
A 2-ended group $H$ is orientable if each element $h\in H$ fixes $\partial H$ pointwise. Otherwise $H$ is non-orientable. If $H$ is orientable, an orientation for $H$ is a choice of any order on the set $\partial H$.
Note that $H$ as above is non-orientable iff for some $h\in H$ the induced homeomorphism of $\partial K$ transposes the points of the doubleton $\partial H$. Assuming now that $H$ is a maximal 2-ended subgroup of a hyperbolic group $K$, if $H$ is orientable then any conjugate of $H$ in $K$ is also orientable, and a similar property holds for non-orientability. Moreover, if $H$ is orientable, any fixed orientation for $H$ induces orientations for all conjugates $gHg^{-1}$. In particular, an orientation for $H$ induces also orders on the boundary sets $\partial(gHg^{-1})$ (which are doubletons). (Here we use the fact that the stabilizer of the doubleton $\partial H\subset\partial K$ coincides with $H$.) As a matter of facts, $H$ is non-orientable iff for some $h\in H$ the left multiplication by $h$ reverses the ends of $H$, so that non-orientability (and orientability) is in fact an intrinsic property of $H$. However, for the purposes of this paper, the corresponding earlier metioned “extrinsic” manifestation of this property is more important.
@@
Orientation data for an abstract factor $(K,\{H_{i}\}_{i\in I})$ is information about orientability and non-orientability of the subgroups $H_{i}$ (and thus of all subgroups from their conjugacy classes) and, for each orientable subgroup $H_{i}$, a choice of an orientation. Such orientation data induces a decoration of the associated space with peripherals, which consists of information about the types ($H_{i}$-types) of the peripherals, and about their orientability, and if a peripheral is orientable, this decoration indicates also one of the orientations of this peripheral, namely this one which is compatible with the fixed orientation of the corresponding 2-ended subgroup.
@@
    Each homeomorphism of the action of $K$ on $\partial K$ preserves any decoration of the associated space with peripherals. More precisely, this means that any such homeomorphism preserves the types of the peripherals, it maps non-orientable peripherals to non-orientable ones, and it maps orientable peripherals to orientable ones, and in the latter case, it preserves the orientations indicated by the decoration.
@@ 8.2. Flexible factors
An abstract factor $(K,\{H_{i}\}_{i\in I})$ is called an abstract flexible factor if the associated space with peripherals is homeomorphic (as a space with peripherals) to the punctured circle $(S^{1})^{\circ}$.
@@
Let $G(v)$ be a flexible factor of the Bowditch JSJ-splitting of a 1-ended hyperbolic group $G$. Then $(G(v),{\mathcal{H}}_{v})$, i.e., the abstract factor associated to $G(v)$ as in Example 8.2, is an abstract flexible factor, in the sense of Definition 8.7. For justification, see Section 5 in [Bow98], where flexible factors are called maximal hanging Fuchsians (MHF), or the paragraph after Definition 5.9 in [KW22].
@@
    Recall that the natural embedding of the punctured circle $(S^{1})^{\circ}$ in the circle $S^{1}$ is uniquely described, up to a homeomorphism of the ambient $S^{1}$, by the following condition: for any peripheral doubleton $P$ of $(S^{1})^{\circ}$ there is a segment $I_{P}$ contained in $S^{1}$ whose set of endpoints coincides with $P$ and whose interior is disjoint with $(S^{1})^{\circ}$. We will call this embedding the natural extension of $(S^{1})^{\circ}$ to a circle. Each segment $I_{P}$ as above will be called the segment induced by $P$. Note that any homeomorphism $h$ of $(S^{1})^{\circ}$ preserving the family of peripherals can be extended in an essentially unique way to a homeomorphism $\bar{h}:S^{1}\to S^{1}$ of the natural extension. Thus, it makes sense to speak of orientation-preserving and orientation-reversing homeomorphisms $h$, by referring to the corresponding property of the induced $\bar{h}$.
@@
Note that an abstract flexible factor $(K,\{H_{i}\}_{i\in I})$ is non-orientable if for some $g\in K$ the induced homeomorphism of $\partial K=(S^{1})^{\circ}$ is orientation-reversing. The next lemma relates orientability properties of an abstract flexible factor $(K,\{H_{i}\}_{i\in I})$ with orientability properties of its peripheral subgroups $H_{i}$. We omit the straightforward proof.
@@
The reader should keep in mind that, given an abstract flexible factor $\mathcal{F}$, orientability of all its peripheral subgroups does not necessarily lead to orientability of the factor. For example, if $K$ is the fundamental group of a compact connected hyperbolic non-orientable surface with non-empty boundary, and if we take as its peripherals the subgroups corresponding to boundary components, then the obtained flexible factor is non-orientable, though all its peripheral subgroups are cyclic, and hence orientable.
@@
Given an orientable abstract flexible factor ${\mathcal{F}}=(K,\{H_{i}\}_{i\in I})$, with a fixed orientation, and an orientable peripheral subgroup $H\in\{H_{i}\}_{i\in I}$, also with a fixed orientation (i.e. an order of the doubleton $P=\partial H$), we say that these orientations are compatible if the orientation induced from $P$ on the induced segment $I_{P}$ in the extended circle $S^{1}$ coincides with the orientation of this segment induced from the orientation of the whole $S^{1}$ which represents the fixed orientation of $\mathcal{F}$.
@@
    Suppose that $\mathcal{F}$ is non-orientable. Consider the extended circle $S^{1}$ for $\partial K=(S^{1})^{\circ}$, and fix one of its orientations. If a subgroup $H_{i}$ is orientable, and if one of its orientations is fixed, consider the induced orientations for all conjugates $gH_{i}g^{-1}$, and the corresponding orientations of all $H_{i}$-peripherals. Then the family of those $H_{i}$-peripherals whose just described orientations are compatible (respectively, incompatible) with the above fixed orientation of $S^{1}$ is non-empty, and the union of this family is dense in $\partial K=(S^{1})^{\circ}$.
@@ 8.3. Abstract virtually free factors and their Whitehead graphs
In this subsection we first recall the concept of a Whitehead graph, as applied to abstract factors that are free. We then extend all of our discussion to abstract virtually free factors. The main result in this subsection, Corollary 8.25, identifies the space with peripherals associated to an abstract virtually free factor with the core of the associated V-tree of Whitehead graphs. This result will play a crucial role in our proof of the implication (3$\Rightarrow$(2) of Theorem A, given in Section 13, and it will be applied to virtually free rigid cluster factors of the reduced Bowditch JSJ splitting of a group $G$.
@@
To each abstract free factor $(F,{\mathcal{H}})$, equipped with a basis $B$ of free generators for $F$, there is associated some finite graph, called the Whitehead graph of the factor, see Section 2E in [Cas16]. We now recall the description of this graph.
@@
Given an abstract free factor ${\mathcal{Q}}=(F,\{H_{i}\})_{1\leq i\leq k}$, let $B=(x_{1},\dots,x_{m})$ be some base set of free generators for $F$. For each $1\leq i\leq k$ fix a generator $h_{i}$ of the cyclic subgroup $H_{i}$, and denote by $w_{i}$ the unique cyclically reduced word over $B\cup B^{-1}$ in the conjugacy class of $h_{i}$. (Here we view $w_{i}$ as a cyclic word, with letters occuring in cyclic order rather than in linear one.) The Whitehead graph for $\mathcal{Q}$ with respect to $B$ is the graph $W=W_{\mathcal{Q}}$ with the vertex set $V_{W_{\mathcal{Q}}}=\{x_{1},\dots,x_{m},x_{1}^{-1},\dots,x_{m}^{-1}\}$ and with one edge connecting distinct vertices $x,y\in V_{W_{\mathcal{Q}}}$ for every occurence of the subword $x^{-1}y$ in the cyclic words $w_{1},\dots,w_{k}$. (We use the convention that if $w_{i}=x$ has length 1 then there is precisely one occurence of $xx$ in $w_{i}$, and no occurence of any other word of length 2; moreover, if $w_{i}=uz$ has length 2, then $w_{i}$ has two subwords of length 2, namely $uz$ and $zu$.)
@@
    Observe that, since the subgroups $H_{i}<F$ are maximal cyclic, their generators are not the powers in $F$, and hence the cyclic words $w_{i}$ are not periodic. For this reason, any two occurences of the same subword in any fixed $w_{i}$ are essentially distinct. This dispels some potential doubts concerning the above definition.
@@
    Note that the Whitehead graph $W_{\mathcal{Q}}$ is obviously a graph with no loop edges, but it may contain multiple edges. $W_{\mathcal{Q}}$ depends not only on the factor $\mathcal{Q}$, but also on the choice of a basis $B$ in the free group $F$. The choices for which the number of edges in corresponding Whitehead graphs is minimal lead to so called minimal Whitehead graphs for the factor $\mathcal{Q}$. As we will show later (see Proposition 9.3), in the situations considered in this paper the minimal Whitehead graphs will always be 2-connected (i.e. finite, connected, nonempty, not equal to a single vertex and cutpoint-free), so in particular they will be essential (i.e. nonempty and with no isolated vertices). Since all our further observations apply to any choice of a minimal Whitehead graph for a factor $\mathcal{Q}$, we will not bother about this choice, and we will speak of the Whitehead graph associated to $\mathcal{Q}$, meaning any of the minimal Whitehead graphs as above.
Any Whitehead graph $W_{\mathcal{Q}}$ which is essential is canonically equipped with some $V$-involution ${\mathbf{a}}_{\mathcal{Q}}$, so that it yields an associated connecting $V$-system ${\mathcal{V}}_{\mathcal{Q}}=(W_{\mathcal{Q}},{\mathbf{a}}_{\mathcal{Q}})$. This $V$-involution consists of the involution $a_{\mathcal{Q}}:V_{W_{\mathcal{Q}}}\to V_{W_{\mathcal{Q}}}$ given by $a_{\mathcal{Q}}(x)=x^{-1}$ for each $x\in V_{W_{\mathcal{Q}}}$ (where we use the convention that $(x_{i}^{-1})^{-1}=x_{i}$), and of the maps $\alpha_{x}:\hbox{Lk}(x)\to\hbox{Lk}(x^{-1})$ described as follows. For any $y\in V_{W_{\mathcal{Q}}}$ and any occurence of $y$ in words $w_{1},\dots,w_{k}$ consider the subwords of the form $xy$ and $yz$, one for each of the two forms, which extend this occurence of $y$, and denote by $\sigma,\tau$ the corresponding occurences of these subwords, respectively. Denote by $e_{\sigma}$ and $e_{\tau}$ the edges of $W$ corresponding to the above occurences of $xy$ and $yz$. Denote also by $[e_{\sigma}]_{y}$ and $[e_{\tau}]_{y^{-1}}$ the points in the links $\hbox{Lk}(y)$ and $\hbox{Lk}(y^{-1})$ induced in those links by the edges $e_{\sigma}$ and $e_{\tau}$, respectively. Finally, put $\alpha_{y}([e_{\sigma}]_{y})=[e_{\tau}]_{y^{-1}}$ and $\alpha_{y^{-1}}([e_{\tau}]_{y^{-1}})=[e_{\sigma}]_{y}$; note that the collection of all assignments as above fully describes all the maps $\alpha_{y}:y\in V_{W_{\mathcal{Q}}}$, and that the conditions required in Definition 7.1 are satisfied. We skip straightforward details of the justification.
The next observation follows directly from Lemma 7.15, in view of the fact that the involution $a_{\mathcal{Q}}$, as given above, is fixed point free.